The accompanying table describes results from groups of 10 births from 10 different sets of parents. The random variable x represents the number of girls among 10 children. Use the range rule of thumb to determine whether 1 girl in 10 births is a significantly low number of girls.

LOADING...

Click the icon to view the table.

x P(x)

0 0.005

1 0.018

2 0.044

3 0.117

4 0.198

5 0.245

6 0.206

7 0.114

8 0.037

9 0.015

10 0.001


1
Expert's answer
2021-02-26T04:54:07-0500

Solution

Range rule of thumb states that the range is approximately four times the standard deviation.





"SD(x) = \sqrt{ \sum {x^2P(x)} - {(\sum {xP(x)}}) ^2}"

"= \sqrt {27.225 - {4.949}^2} =1.652997"

"Range = 4 SD(x)"

"=4(1.652997)=6.611988"


But

"Range = max. \space value - min. \space value"

"Max. = \bar{x} + 2SD(x)""=4.949+2(1.652997) =8.254994"

"Min = \bar {x} - 2SD(x)""=4.949-2(1.652997) =1.643"

Since

"Maximum \approx 9 \space"and "Minimum \approx 2"


1 girl in 10 births is a significantly low number of girls